A guide dog training center took half of the puppies at the dog pound. Steve took half of the remaining. Tanya took the 5 that were left. How many puppies were there to begin with?

Answer by TimothyLamb(4379)

g = D/2
s = (D - D/2)/2
s = (D/2 - D/4)
s = (2D/4 - D/4) = D/4
D - g - s = 5
D - D/2 - D/4 = 5
---
D - D/2 - D/4 = 5
4D/4 - 2D/4 - D/4 = 5
D/4 = 5
D = 20
guide = 10
steve = 5
tanya = 5
---
total pups: 20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php